OpenStudy (anonymous):

Solve for x: 1.(tanx/cotx)(-secx/cosx)=2/cosx Verify 2.(cot^2x+1)(sin^2x-1)=-cot2X 3. (sinx+1)/(sinx)=1+cscx

OpenStudy (anonymous):

the last one is the easiest

OpenStudy (anonymous):

\[\frac{\sin(x)+1}{\sin(x)}=\frac{\sin(x)}{\sin(x)}+\frac{1}{\sin(x)}=1+\csc(x)\]

OpenStudy (anonymous):

first one start with the fact that \(\frac{\tan(x)}{\cot(x)}=1\)

OpenStudy (anonymous):

also \(\frac{\sec(x)}{\cos(x)}=1\) so the left hand side is \(-1\)

OpenStudy (anonymous):

you get \[-1=\frac{2}{\cos(x)}\] so \(\cos(x)=-2\) which is not possible